# Odd Dipole Screening in Radial Inflation

### v1

@@ I Introduction
In a series of recent papers the phenomenon of screening in the mechanical response of amorphous solids to non-uniform stresses or strains was examined theoretically[1, 2, 3, 4, 5], experimentally [6, 7] and simulationally [8, 9, 10]. The phenomenon at question deals with an amorphous solid that is at mechanical equilibrium (i.e. with vanishing net force on each constituent particle), that is subjected to strain, leading to loss of mechanical equilibrium. Upon relaxation back to mechanical equilibrium, the positions of constituent particles change, until a new configuration is attained in which the net force on each particle vanishes. We are interested in the displacement field ${\bm{d}}({\bm{r}})$ which is the difference between the final and initial in positions of the centers of mass of each particle.
For a classical elastic solid the equation obeyed by displacement field ${\bm{d}}$ reads [11]:
@@
Here $\lambda$ and $\mu$ are the classical Lame’ coefficients. For an amorphous solids things are different. First, generically any strain on amorphous solids results in plastic deformation [12, 13]. Typical plastic events are quadrupolar in nature [14, 15], called sometime “Eshelby inclusions" due to the resemblance to the Eshelby solution of the displacement field caused by forcing a circle to change to an ellipse within an elastic material [16]. When the density of quadrupolar events is finite, one refers to the resulting quadrupolar field as $Q^{\alpha\beta}({\bm{r}})$. It was shown the creation of such a field leads to a renormalization of the elastic moduli [1, 2]. Further, when the quadrupolar field of such events is non-uniform, it was explained that gradients of the field act as effective dipoles,
@@
When these are present, the equation satisfied by the displacement field change drastically, reading [1, 2]
@@
This constitutive relations means that the dipole and the displacement fields are expected to be co-linear and opposite in direction. This prediction was not accurately corroborated in simulations, an angle was distinctly existing between the prediction and the measured direction of the dipole field.
The resolution of this issue was offered in Ref. [7]. It was argued that in systems that do not conserve energy along a closed loop of straining, the assumption of diagonal tensor ${\bm{\Gamma}}$ is not tenable. Rather, an odd tensor is expected, of the form
${\bm{\Gamma}}=\begin{bmatrix}\kappa_{e}^{2}&-\kappa_{o}^{2}\\ \kappa_{o}^{2}&\kappa_{e}^{2}\end{bmatrix}\ .$ (6)
The presence of this odd tensor in the theory has led to the term “odd dipole screening". The existence of the anti-symmetric tensor ${\bm{\Gamma}}$ leads to breaking Chiral symmetry. Even for purely radial inflation as discussed below, the resulting displacement field exhibits non-trivial transverse (angular) response, as will be shown explicitly in the next Section. Denoting the angle formed between the (inverse) direction of the displacement field and the dipole field as $\Phi$, one predicts that
$\tan\Phi=\frac{\kappa_{o}^{2}}{\kappa_{e}^{2}}\ .$ (7)
Tests of this prediction are offered below.
@@ II.1 Solving for the displacement field
We then evaluate the screening term as
@@
Now equation (3) can be decomposed into following coupled differential equations in $r,$ and $\theta$ :
@@
Figure 2: Panel (a) and (b): Typical solutions of $a_{0}(r)$ and $b_{0}(r)$ for a fixed $\kappa_{\rm o}$=0.041 and for different values of $\kappa_{\rm e}=0.0204,0.0612,0.102,0.1429$. Panel (c): Comparison of $a_{0}$ and $b_{0}$ for $\kappa_{\rm o}=0.041$ and $\kappa_{\rm e}=0.204$.
Figure 3: Panel (a) and (b): Typical solutions of $a_{0}(r)$ and $b_{0}(r)$ for a fixed $\kappa_{\rm e}$=0.041 and for different values of $\kappa_{\rm o}=0.0204,0.0612,0.102,0.1429$. Panel (c): Comparison of $a_{0}$ and $b_{0}$ for $\kappa_{\rm o}=0.204$ and $\kappa_{\rm e}=0.041$.
@@
$d_{r}(r,\theta)=a_{0}(r)+\sum_{n=1}^{\infty}\left[a_{n}(r)\cos(n\theta)+b_{n}(r)\sin(n\theta)\right],$
$d_{\theta}(r,\theta)=b_{0}(r)+\sum_{n=1}^{\infty}\left[c_{n}(r)\cos(n\theta)+d_{n}(r)\sin(n\theta)\right].$ (12)
@@
$d_{r}(r,\theta)=a_{0}(r)+a_{1}(r)\cos(\theta)+b_{1}(r)\sin(\theta),$
$d_{\theta}(r,\theta)=b_{0}(r)+c_{1}(r)\cos(\theta)+d_{1}(r)\sin(\theta).$ (13)
@@
$+\left[\left[r^{2}a_{1}^{\prime\prime}(r)+ra_{1}^{\prime}(r)\right]+\frac{(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)}{(\tilde{\lambda}+2)}a_{1}(r)-\frac{\kappa_{o}^{2}r^{2}}{(\tilde{\lambda}+2)}c_{1}(r)+\frac{(\tilde{\lambda}+1)r}{(\tilde{\lambda}+2)}d_{1}^{\prime}(r)-\frac{(\tilde{\lambda}+3)}{(\tilde{\lambda}+2)}d_{1}(r)\right]\cos(\theta)$
$+\left[\left[r^{2}b_{1}^{\prime\prime}(r)+rb_{1}^{\prime}(r)\right]+\frac{(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)}{(\tilde{\lambda}+2)}b_{1}(r)-\frac{\kappa_{o}^{2}r^{2}}{(\tilde{\lambda}+2)}d_{1}(r)-\frac{(\tilde{\lambda}+1)r}{(\tilde{\lambda}+2)}c_{1}^{\prime}(r)+\frac{(\tilde{\lambda}+3)}{(\tilde{\lambda}+2)}c_{1}(r)\right]\sin(\theta)=0,$ (14)
@@
$+\left[\left[r^{2}c_{1}^{\prime\prime}(r)+rc_{1}^{\prime}(r)\right]+(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)c_{1}(r)+\kappa_{o}^{2}r^{2}a_{1}(r)+(\tilde{\lambda}+1)rb_{1}^{\prime}(r)+(\tilde{\lambda}+3)b_{1}(r)\right]\cos(\theta)$
$+\left[\left[r^{2}d_{1}^{\prime\prime}(r)+rd_{1}^{\prime}(r)\right]+(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)d_{1}(r)+\kappa_{o}^{2}r^{2}b_{1}(r)-(\tilde{\lambda}+1)ra_{1}^{\prime}(r)-(\tilde{\lambda}+3)a_{1}(r)\right]\sin(\theta)=0.$ (15)
@@ II.2 Analytic solutions
The equations for $a_{0}$ and $b_{0}$ can be solved analytically, as is demonstrated in Appendix A. The final solutions are shown in Eqs. (55) and (56). Typical solutions are shown in Figs. 2 and 3.
@@ II.3 Comparison with simulations
The description of the numerical procedures for the preparation of our amorphous solids for the inflation of the inner disk is offered in Appendix B.
@@
Typical comparisons of numerical and exact solution are shown in Fig. 5.
Figure 5: Comparisons of angle averaged radial ($a_{0}(r)$) and transverse ($b_{0}(r)$) displacement field with that of analytical solutions (solid line). Upper figure: $r_{in}=1.25$, $r_{out}=80$, $\mu=3.43$, $\tilde{\lambda}=68.9$, $k_{e}=0.209$ and $k_{o}=0.206$. Here $a_{0}(r_{in})=0.16$ and $b_{0}(r_{in})=-0.2$. Middle figure: $r_{in}=1.3$, $r_{out}=80$, $\mu=5.31$, $\tilde{\lambda}=41.3$, $k_{e}=0.099$ and $k_{o}=0.149$. Here $a_{0}(r_{in})=0.18$ and $b_{0}(r_{in})=-0.02$. Lower figure: $r_{in}=1.25$, $r_{out}=80$, $\mu=3.73$, $\tilde{\lambda}=39$, $k_{e}=0.165$ and $k_{o}=0.155$. Here $a_{0}(r_{in})=0.19$ and $b_{0}(r_{in})=-0.08$.
We note that the analytic solutions depend on the actual elastic moduli $\mu$ and $\lambda$. We have computed these for the actual realizations after equilibration, from the microscopic theory that is described for example in Ref. [17]. In contrast, the screening parameters $\kappa_{e}$ and $\kappa_{o}$ were fitted to the data, and were the only free parameters. We find the agreement between data and theory quite satisfactory, despite some slight deviation mainly in $b_{0}(r)$ at higher values of $r$. We therefore turn now to an even more sensitive test for the fit of the screening parameters, i.e. the angle $\Phi$ defined in Eq. (7).
@@ III.1 Extracting the Quadrupolar and Dipole Fields from displacement data
To accomplish our aim we must extract first the quadrupolar field associated with the displacement field. Since a quadrupole is represented by the traceless part of the non-affine strain tensor, so we first compute the total strain tensor. We follow the procedure advised in Ref. [18], using the discrete deformation gradient $F$ defined as,
@@
To calculate $F$ in the atomistic simulation we measure the relative displacements between two particles. Before we proceed, we would like to mention that the particle positions have been mapped to grid points. Therefore, in all our subsequent discussion particle labels and grid points are synonyms. Let $\Delta X^{ij}=X^{j}-X^{i}$, and $\Delta x^{ij}=x^{j}-x^{i}$ be the relative displacements in $\Omega_{0}$ and $\Omega_{D}$ respectively between the particles $i$ and $j$. The transformation from configuration $\Omega_{0}$ to $\Omega_{D}$ is defined by deformation gradient $F$, which linearly maps the relative position vector $\bf{\Delta X^{ij}}$ to $\bf{\Delta x^{ij}}$ as follows,
@@
where $F^{i}$ is the deformation gradient measured at the particle $i$. Note that, instead of $F$, we use a local deformation gradient $\hat{F}$, which is given by [18],
@@
Here $w_{j}$ is the weight function assigned to particle $j$ depending upon its relative distance with the central atom $i$, and $L$ denotes the total neighbors around particle $i$ [18].
@@
Here $\varepsilon_{\alpha\beta}^{a}=0.5(\nabla_{\alpha}d_{\beta}+\nabla_{\beta}d_{\alpha})$, and $\alpha,\beta=1,2$ which are same as $(x,y)$. Note that $\varepsilon_{\alpha\beta}^{q}$ is calculated at each particle label or each grid point. Here, $d$ is the quasi-elastic displacement field in purely radial inflation which is given by [1]:
$d(r)=d_{o}\left[\frac{r^{2}-r_{out}^{2}}{r_{in}^{2}-r_{out}^{2}}\right]\left(\frac{r_{in}}{r}\right).$ (22)
Figure 6: Heat map of the quadrupolar field in typical radial inflation. The dense regions indicate high values of $Q$, and less dense regions correspond to the low values of $Q$ as obtained from Eq. (24). The arrows represent the principal axis or the directions of the $Q$ field, cf. Eq. (26). The parameters here are $r_{in}=1.25$, $r_{out}=80$, $\mu=3.43$, $\tilde{\lambda}=68.9$, $k_{e}=0.209$ and $k_{o}=0.206$. Here $a_{0}(r_{in})=0.16$ and $b_{0}(r_{in})=-0.2$. In the top panel we represent the full quadrupolar field, whereas in the bottom panel we show the magnified view of the most dense region.
The non-affine strain $\varepsilon^{q}$ can be decomposed into its trace and its traceless components [10]:
$\varepsilon^{q}$ $=\frac{1}{2}\begin{bmatrix}\varepsilon_{11}^{q}+\varepsilon_{22}^{q}&0\\ 0&\varepsilon_{11}^{q}+\varepsilon_{22}^{q}\end{bmatrix}+Q\begin{bmatrix}\varepsilon_{11}^{ts}&\varepsilon_{12}^{ts}\\ \varepsilon_{21}^{ts}&\varepsilon_{22}^{ts}\end{bmatrix},$
$=0.5\times mI+Q\varepsilon^{ts}.$ (23)
Here $I$ is the identity tensor, $m=(\varepsilon_{11}^{q}+\varepsilon_{22}^{q}$) is the trace of the tensor $\varepsilon^{q}$, and $\varepsilon^{ts}$ is traceless symmetric tensor representing a quadrupole, where $\varepsilon_{22}^{ts}=-\varepsilon_{11}^{ts}$. The quantity $Q$ is the quadrupolar charge associated with the quadrupolar tensor and is defined as,
$Q^{2}=(\varepsilon_{11}^{ts})^{2}+(\varepsilon_{22}^{ts})^{2}.$ (24)
Eq. (III.1) can also be written in the following form,
@@
$\Theta$ defines the principal axis of the quadrupolar field. We show a typical quadrupolar field in Fig. 6.
@@
Figure 7: Heat map of the dipole field after a typical radial inflation, obtained by taking the gradient of the quadrupolar field ${\mathcal{P}}^{\alpha}\equiv\partial_{\beta}Q^{\alpha\beta}$. The dense regions indicate high values of ${\mathcal{P}}$, and less dense regions correspond to the low values of dipole field ${\mathcal{P}}$. The arrows represent the local direction of the dipole field. The parameters here are are $r_{in}=1.25$, $r_{out}=80$, $\mu=3.43$, $\tilde{\lambda}=68.9$, $k_{e}=0.209$ and $k_{o}=0.206$. Here $a_{0}(r_{in})=0.16$ and $b_{0}(r_{in})=-0.2$. In the top panel we represent the full dipole field, whereas in the bottom panel we show the magnified view of the most dense region.
@@
Having the dipole field we compute the angle between the inverse displacement and the dipole field, which is predicted to be the angle $\Phi$ of Eq. (7). For the present example, we found from simulations an angle of $46.48^{o}\pm 0.55^{o}$ as compared to an angle of $44.11^{o}$ from the direct calculation of Eq. (7). This close correspondence is typical and similar results were obtained for other simulations of radial inflation.

### v3

@@ I Introduction
In a series of recent papers the phenomenon of screening in the mechanical response of amorphous solids to non-uniform stresses or strains was examined theoreticallyLemaître et al. (2021); Charan et al. (2023); Kumar and Procaccia (2024); Jin et al. (2024); Hentschel et al. (2024), experimentally Mondal et al. (2022); Cohen et al. (2024) and simulationally Bhowmik et al. (2022); Kumar et al. (2022); Mondal et al. (2023). The phenomenon at question deals with an amorphous solid that is at mechanical equilibrium (i.e. with vanishing net force on each constituent particle), that is subjected to strain, leading to loss of mechanical equilibrium. Upon relaxation back to mechanical equilibrium, the positions of constituent particles change, until a new configuration is attained in which the net force on each particle vanishes. We are interested in the displacement field ${\bm{d}}({\bm{r}})$ which is the difference between the final and initial in positions of the centers of mass of each particle.
For a classical elastic solid the equation obeyed by displacement field ${\bm{d}}$ reads Landau and Lifshitz (1959):
@@
Here $\lambda$ and $\mu$ are the classical Lame’ coefficients. For an amorphous solids things are different. First, generically any strain on amorphous solids results in plastic deformation Karmakar et al. (2010); Hentschel et al. (2011). Typical plastic events are quadrupolar in nature Malandro and Lacks (1999); Maloney and Lemaître (2006), called sometime “Eshelby inclusions" due to the resemblance to the Eshelby solution of the displacement field caused by forcing a circle to change to an ellipse within an elastic material Eshelby (1957). When the density of quadrupolar events is finite, one refers to the resulting quadrupolar field as $Q^{\alpha\beta}({\bm{r}})$. It was shown the creation of such a field leads to a renormalization of the elastic moduli Lemaître et al. (2021); Charan et al. (2023). Further, when the quadrupolar field of such events is non-uniform, it was explained that gradients of the field act as effective dipoles,
@@
When these are present, the equation satisfied by the displacement field change drastically, reading Lemaître et al. (2021); Charan et al. (2023)
@@
This constitutive relations means that the dipole and the displacement fields are expected to be co-linear and opposite in direction. This prediction was not accurately corroborated in simulations, an angle was distinctly existing between the prediction and the measured direction of the dipole field Mondal et al. (2023).
The resolution of this issue was offered in Ref. Cohen et al. (2024). Eq. (3) was derived Lemaître et al. (2021); Charan et al. (2023) from a potential energy, which implies a symmetric tensor $\Gamma^{\alpha\beta}=\Gamma^{\beta\alpha}$. However, in typical amorphous solids energy is not conserved due to plastic rearrangements. Straining along a closed loop in strain space is likely to bring the system to a new state rather than the initial one. This suggests that a diagonal tensor ${\bm{\Gamma}}$ cannot tell the whole story. To remove the constraint of energy conservation, but keeping a successful screening theory, we allow ${\bm{\Gamma}}$ to contain an anti-symmetric screening parameter, of the form
${\bm{\Gamma}}=\begin{bmatrix}\kappa_{e}^{2}&\mp\kappa_{o}^{2}\\ \pm\kappa_{o}^{2}&\kappa_{e}^{2}\end{bmatrix}\ .$ (6)
The $\mp$ sign of $\kappa_{0}^{2}$ means that it can be negative or positive, with an anti-symmetric counterpart ±$\pm\kappa_{0}^{2}$. The presence of this odd tensor in the theory has led to the term “odd dipole screening". The existence of the anti-symmetric tensor ${\bm{\Gamma}}$ leads to breaking Chiral symmetry. Even for purely radial inflation as discussed below, the resulting displacement field exhibits non-trivial transverse (angular) response, as will be shown explicitly in the next Section. Denoting the angle formed between the (inverse) direction of the displacement field and the dipole field as $\Phi$, one predicts that
$\tan\Phi=\pm\frac{\kappa_{o}^{2}}{\kappa_{e}^{2}}\ .$ (7)
The angle $\Phi$ can be positive or negative, depending on the sign of $\kappa_{o}^{2}$, but is expected to be the same for all the angles between the dipole and displacement fields in a given realization Cohen et al. (2024). Tests of this prediction are offered below.
@@ II.1 Solving for the displacement field
For concreteness we will present the solution for the case of $\kappa_{o}^{2}$ positive, and a parallel analysis can be easily done for a negative $\kappa_{o}^{2}$. We then evaluate the screening term as
@@
Now equation (3) can be decomposed into following coupled differential equations in $r,$ and $\theta$ (with prime and double prime standing for first or second space derivatives):
@@
Figure 2: Panel (a) and (b): Typical solutions of $a_{0}(r)$ and $b_{0}(r)$ for a fixed $\kappa_{\rm o}$=0.041 and for different values of $\kappa_{\rm e}=0.0204,0.0612,0.102,0.1429$. Panel (c): Comparison of $a_{0}$ and $b_{0}$ for $\kappa_{\rm o}=0.041$ and $\kappa_{\rm e}=0.204$. In all these cases the boundary conditions are $a_{0}(r_{\rm in})=0.1$, $a_{0}(r_{\rm out})=0$, $b_{0}(r_{\rm in})=0$,$b_{0}(r_{\rm out})=0$.
Figure 3: Panel (a) and (b): Typical solutions of $a_{0}(r)$ and $b_{0}(r)$ for a fixed $\kappa_{\rm e}$=0.041 and for different values of $\kappa_{\rm o}=0.0204,0.0612,0.102,0.1429$. Panel (c): Comparison of $a_{0}$ and $b_{0}$ for $\kappa_{\rm o}=0.204$ and $\kappa_{\rm e}=0.041$. The boundary conditions are the same as in Figs. 2
@@
$d_{r}(r,\theta)=a_{0}(r)+\sum_{n=1}^{\infty}\left[a_{n}(r)\cos(n\theta)+c_{n}(r)\sin(n\theta)\right],$
$d_{\theta}(r,\theta)=b_{0}(r)+\sum_{n=1}^{\infty}\left[g_{n}(r)\cos(n\theta)+e_{n}(r)\sin(n\theta)\right].$ (12)
@@
$d_{r}(r,\theta)=a_{0}(r)+a_{1}(r)\cos(\theta)+c_{1}(r)\sin(\theta),$
$d_{\theta}(r,\theta)=b_{0}(r)+g_{1}(r)\cos(\theta)+e_{1}(r)\sin(\theta).$ (13)
@@
$+\left[\left[r^{2}a_{1}^{\prime\prime}(r)+ra_{1}^{\prime}(r)\right]+\frac{(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)}{(\tilde{\lambda}+2)}a_{1}(r)-\frac{\kappa_{o}^{2}r^{2}}{(\tilde{\lambda}+2)}g_{1}(r)+\frac{(\tilde{\lambda}+1)r}{(\tilde{\lambda}+2)}e_{1}^{\prime}(r)-\frac{(\tilde{\lambda}+3)}{(\tilde{\lambda}+2)}e_{1}(r)\right]\cos(\theta)$
$+\left[\left[r^{2}c_{1}^{\prime\prime}(r)+rc_{1}^{\prime}(r)\right]+\frac{(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)}{(\tilde{\lambda}+2)}c_{1}(r)-\frac{\kappa_{o}^{2}r^{2}}{(\tilde{\lambda}+2)}e_{1}(r)-\frac{(\tilde{\lambda}+1)r}{(\tilde{\lambda}+2)}g_{1}^{\prime}(r)+\frac{(\tilde{\lambda}+3)}{(\tilde{\lambda}+2)}g_{1}(r)\right]\sin(\theta)=0,$ (14)
@@
$+\left[\left[r^{2}g_{1}^{\prime\prime}(r)+rg_{1}^{\prime}(r)\right]+(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)g_{1}(r)+\kappa_{o}^{2}r^{2}a_{1}(r)+(\tilde{\lambda}+1)rc_{1}^{\prime}(r)+(\tilde{\lambda}+3)c_{1}(r)\right]\cos(\theta)$
$+\left[\left[r^{2}e_{1}^{\prime\prime}(r)+re_{1}^{\prime}(r)\right]+(\kappa_{e}^{2}r^{2}-\tilde{\lambda}-3)e_{1}(r)+\kappa_{o}^{2}r^{2}c_{1}(r)-(\tilde{\lambda}+1)ra_{1}^{\prime}(r)-(\tilde{\lambda}+3)a_{1}(r)\right]\sin(\theta)=0.$ (15)
@@ II.2 Analytic solutions
The equations for $a_{0}$ and $b_{0}$ can be solved analytically, as is demonstrated in Appendix A. The final solutions are shown in Eqs. (56) and (57). Typical solutions are shown in Figs. 2 and 3. Here $\tilde{\lambda}=9.17$. We note that in the limit $\tilde{\lambda}\to\infty$ the analytic solutions converge the the purely elastic solution.
@@ II.3 Comparison with simulations
The description of the numerical procedures for the preparation of our amorphous solids for the inflation of the inner disk is offered in Appendix B. In all the simulations below the inflation magnitude $d_{0}$ is between 10% to 20% of $r_{\rm in}$.
@@
Typical comparisons of numerical and exact solution are shown in Fig. 5. Here are three independent realizations, with different material parameters, and different geometries.
Figure 5: Comparisons of angle averaged radial ($a_{0}(r)$) and transverse ($b_{0}(r)$) displacement field with that of analytical solutions (solid line). Upper figure: $r_{in}=1.25$, $r_{out}=80$, $\tilde{\lambda}=20.09$, $\kappa_{e}=0.209$ and $\kappa_{o}=0.206$. Here $a_{0}(r_{in})=0.16$ and $b_{0}(r_{in})=-0.2$. Middle figure: $r_{in}=1.3$, $r_{out}=80$, $\tilde{\lambda}=7.9$, $\kappa_{e}=0.099$ and $\kappa_{o}=0.149$. Here $a_{0}(r_{in})=0.18$ and $b_{0}(r_{in})=-0.02$. Lower figure: $r_{in}=1.25$, $r_{out}=80$, $\tilde{\lambda}=10.46$, $\kappa_{e}=0.165$ and $\kappa_{o}=0.155$. Here $a_{0}(r_{in})=0.19$ and $b_{0}(r_{in})=-0.08$.
We note that the analytic solutions depend on the actual elastic moduli $\mu$ and $\lambda$. We have computed these for the actual realizations after equilibration, from the microscopic theory that is described for example in Ref. Lutsko (1989). In contrast, the screening parameters $\kappa_{e}$ and $\kappa_{o}$ were fitted to the data, and were the only free parameters. We find the agreement between data and theory quite satisfactory, despite some slight deviation mainly in $b_{0}(r)$ at higher values of $r$. We therefore turn now to an even more sensitive test for the fit of the screening parameters, i.e. the angle $\Phi$ defined in Eq. (7).
@@ III.1 Extracting the Quadrupolar and Dipole Fields from displacement data
To accomplish our aim we must extract first the quadrupolar field associated with the displacement field. Since a quadrupole is represented by the traceless part of the non-affine strain tensor, so we first compute the total strain tensor. We follow the procedure advised in Ref. Gullett et al. (2007), using the discrete deformation gradient $F$ defined as,
@@
To calculate $F$ in the atomistic simulation we measure the relative displacements between two particles. Before we proceed, we would like to mention that the particle positions have been mapped to grid points. To this aim we have used the Matlab library that uses cubic spline and Linear Interpolation standard methods. Therefore, in all our subsequent discussion particle labels and grid points are synonyms. Let $\Delta X^{ij}=X^{j}-X^{i}$, and $\Delta x^{ij}=x^{j}-x^{i}$ be the relative displacements in $\Omega_{0}$ and $\Omega_{D}$ respectively between the particles $i$ and $j$. The transformation from configuration $\Omega_{0}$ to $\Omega_{D}$ is defined by deformation gradient $F$, which linearly maps the relative position vector $\bf{\Delta X^{ij}}$ to $\bf{\Delta x^{ij}}$ as follows,
@@
where $F^{i}$ is the deformation gradient measured at the particle $i$. Note that, instead of $F$, we use a local deformation gradient $\hat{F}$, which is given by Gullett et al. (2007),
@@
Here $w_{j}$ is the weight function assigned to particle $j$ depending upon its relative distance with the central atom $i$, and $L$ denotes the total neighbors around particle $i$ Gullett et al. (2007).
@@
Here $\varepsilon_{\alpha\beta}^{a}=0.5(\nabla_{\alpha}d^{a}_{\beta}+\nabla_{\beta}d^{a}_{\alpha})$, and $\alpha,\beta=1,2$ which are same as $(x,y)$. Note that $\varepsilon_{\alpha\beta}^{q}$ is calculated at each particle label or each grid point. Here, $d^{a}$ is the quasi-elastic displacement field in purely radial inflation which is given by Lemaître et al. (2021):
$d^{a}(r)=d_{o}\left[\frac{r^{2}-r_{out}^{2}}{r_{in}^{2}-r_{out}^{2}}\right]\left(\frac{r_{in}}{r}\right).$ (22)
Figure 6: Heat map of the quadrupolar field in typical radial inflation. The dense regions indicate high values of $Q$, and less dense regions correspond to the low values of $Q$ as obtained from Eq. (25). The arrows represent the principal axis or the directions of the $Q$ field, cf. Eq. (26). The parameters here correspond to the first pair of panels in Fig. 5. In the top panel we represent the full quadrupolar field, whereas in the bottom panel we show the magnified view of the most dense region.
The non-affine strain $\varepsilon^{q}$ can be decomposed into its trace and its traceless components Mondal et al. (2023):
$\varepsilon^{q}$ $=\frac{1}{2}\begin{bmatrix}\varepsilon_{11}^{q}+\varepsilon_{22}^{q}&0\\ 0&\varepsilon_{11}^{q}+\varepsilon_{22}^{q}\end{bmatrix}+\begin{bmatrix}\varepsilon_{11}^{ts}&\varepsilon_{12}^{ts}\\ \varepsilon_{21}^{ts}&\varepsilon_{22}^{ts}\end{bmatrix}\ .$ (23)
The traceless component has a quadrupolar structure, and Eq. (23) can also be written in the following form,
@@
The quantity $Q$ is the quadrupolar charge associated with the quadrupolar tensor and is defined as,
$Q^{2}=(\varepsilon_{11}^{ts})^{2}+(\varepsilon_{12}^{ts})^{2}.$ (25)
@@
$\Theta$ defines the principal axis of the quadrupolar field. We show a typical quadrupolar field in Fig. 6. The data used to generate this and the next figure correspond to the top system in Fig. 5.
@@
Figure 7: Heat map of the dipole field after a typical radial inflation, obtained by taking the gradient of the quadrupolar field ${\mathcal{P}}^{\alpha}\equiv\partial_{\beta}Q^{\alpha\beta}$. The dense regions indicate high values of ${\mathcal{P}}$, and less dense regions correspond to the low values of dipole field ${\mathcal{P}}$. The arrows represent the local direction of the dipole field. The parameters here correspond to the first pair of panels in Fig. 5. In the top panel we represent the full dipole field, whereas in the bottom panel we show the magnified view of the most dense region.
@@
Having the dipole field we compute the angle between the inverse displacement and the dipole field, which is predicted to be the angle $\Phi$ of Eq. (7). In Fig. 8 we show simultaneously the dipole and displacement fields, in two colors, from which we can determine the angles between them. For the present example, we found from simulations an average angle of $46.48^{o}\pm 0.55^{o}$ as compared to an angle of $44.11^{o}$ from the direct calculation of Eq. (7). This close correspondence is typical and similar results were obtained for other simulations of radial inflation.